OpenStudy (rootbeer003):

Which set if lengths could represent the sides of a triangle? A) 6,7,14 B)3,7,9 C)4,6,10 D)7,8,18

Vocaloid (vocaloid):

any ideas?

OpenStudy (rootbeer003):

I've been stuck on this forever but my idea was b

OpenStudy (sshayer):

sum of any two sides > third side.

OpenStudy (rootbeer003):

Ugh idk

OpenStudy (rootbeer003):

Wait c?

OpenStudy (mertsj):

How could it be C when 4+6 is not bigger than 10?

OpenStudy (rootbeer003):

Then b cuz it's 12

OpenStudy (mertsj):

3,7,9 3+7>9 3+9>7 7+9>3 So yes, B works.

OpenStudy (sshayer):

3+7=10>9 3+9=12>7 7+9=16>3